Balance Sheet

Income Statement
Cash $ 72,000 Accts & Notes Pay. $ 432,000
A/R 439,000 Accruals 170,000
Inventories 894,000 Total current liabilities 602,000
Total current assets 1,405,000
Land and bldg 238,000 Long-term debt 404,290
Machinery 132,000 Common stock 575,000
Other F.A. 61,000 Retained earnings (RE) 254,710
Total assets $ 1,836,000 Total liab. & equity $ 1,836,000
RE last year 146,302

INCOME STATEMENT

Sales $ 4,290,000
Cost of G.S.      3,580,000
Gross operating profit         710,000
Adm. & sales exp.         236,320
Depreciation         159,000
Misc. expenses         134,000
Earnings before taxes         180,680
Taxes (40%)           72,272
Net income $     108,408
P/E ratio                 5.0
No. of shares           23,000
Cash dividend $           0.95
Tax rate 40%

Calculate the ratios. Please show your work.

Quick
Current
Inv. turn.
DSO
FA turnover
TA turnover
ROA
ROE
TD/TA
PM
EPS
Stock Price
P/E ratio
M/B
Answer #1
Quick or acid test ratio = (current assets-inventory)/current liablities
Quick or acid test ratio = (1405000-894000)/602000
Quick or acid test ratio = 0.85
Current ratio = current assets/current liablities
Current ratio = 1405000/602000
Current ratio = 2.33
Inventory turnover = COGS/inventory
Inventory turnover = 3580000/894000
Inventory turnover = 4
Receivables turnover = Credit sales/receivables
Receivables turnover = 4290000/439000
Receivables turnover = 9.77
days of sales outstanding = number of days in a year/receivables turnover
days of sales outstanding = 365/9.77
days of sales outstanding = 37.36
Please ask remaining parts seperately,